The radius of the circle having centre at (2,1) whose one chord is the diameter of the circle x2 +y2 -2x-6y+6=0

Grade:12

1 Answers

Rup Chitrak
12 years ago

x2+y2-2x-6y+6=0

or, (x2-2x+1)+(y2-6y+9)-4=0

or, (x-1)2+(y-3)2=22

thus the radius of the smaller given circle is 2 and its centre is (1,3).

distance b/w the two centres = (1+4)1/2 = 51/2

thus the radius of required circle = (5 + 4)1/2 = 3.